OpenStudy (anonymous):

Find dy/dx of sqrt(x+y) + sqrt(xy) = 6

OpenStudy (anonymous):

=\[sqrt{xy}\/\(sqrt{x+y}\=sqrt{xy}\)\]

OpenStudy (anonymous):

if this is the ans..i will explain...let me know..

OpenStudy (anonymous):

what are the red dashes for? and sorry i do not know the answer of this problem.

OpenStudy (anonymous):

=\[sqrt{xy}\/\(sqrt{x+y}\+sqrt{xy}\)\] red was f0r eqstn editor...it didint worked...

OpenStudy (anonymous):

uh maybe type in normal horizontal equation please :)

OpenStudy (anonymous):

\[\frac{d}{dx}[\sqrt{x+y} + \sqrt{xy}] = 0\]\[\implies \frac{d}{dx}[\sqrt{x+y}] + \frac{d}{dx}[\sqrt{xy}] = 0\]By the chain rule:\[\implies \frac{1}{2\sqrt{x+y}}\cdot \frac{d}{dx}[x+y] + \frac{1}{2\sqrt{xy}}\cdot \frac{d}{dx}[xy] = 0\]\[\implies \frac{1}{2\sqrt{x+y}}(1 + \frac{dy}{dx}) + \frac{1}{2\sqrt{xy}}\cdot (x\frac{dy}{dx} + y) = 0\] And continue from there to solve for \(\frac{dy}{dx}\)

OpenStudy (anonymous):

I gotta run to the store, but if you get stuck, lemme know.

OpenStudy (anonymous):

you refer to dy/dx is just like y' right? 'cause i used y' it's cleaner haha. but that's an awesome and very clean looking solution. thanks a lot. +1